# Deeper Insights into Deep Graph Convolutional Networks: Stability and Generalization

### v1

@@ A.3 Proof of Lemma 3 and Lemma 4
A.3.1 Proof of Lemma 3
To show (15), note that
$=$ $\|\frac{\partial\ell(\hat{y},y_{t})}{\partial\hat{y}}\nabla_{\mathbf{w}}f(\mathbf{x}|\theta_{t-1})-\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime}_{t-1})\|_{F}$
$\leq$ $\|\Big{(}\frac{\partial\ell(\hat{y},y_{t})}{\partial\hat{y}}-\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}\Big{)}\nabla_{\mathbf{w}}f(\mathbf{x}|\theta_{t-1})+\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}\Big{(}\nabla_{\mathbf{w}}f(\mathbf{x}|\theta_{t-1})-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime}_{t-1})\Big{)}\|_{F}$
$\leq$ $|\frac{\partial\ell(\hat{y},y_{t})}{\partial\hat{y}}-\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}|\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta_{t-1})\|_{F}+|\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}|\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta_{t-1})-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime}_{t-1})\|_{F}$
$\leq$ $\upsilon_{\ell}|f(\mathbf{x}|\theta_{t-1})-f(\mathbf{x}|\theta^{\prime}_{t-1})|\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta_{t-1})\|_{F}+\alpha_{\ell}\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta_{t-1})-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime}_{t-1})\|_{F},$
where $\hat{y}=f(\mathbf{x}|\theta_{t-1})$ and $\hat{y}^{\prime}=f(\mathbf{x}|\theta^{\prime}_{t-1})$. Then, according to (25), (27) and (28), we have
@@ A.3.1 Proof of Lemma 3
$\leq$ $\upsilon_{\ell}|f(\mathbf{x}|\theta_{t-1})-f(\mathbf{x}|\theta^{\prime}_{t-1})|\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta_{t-1})\|_{F}+\alpha_{\ell}\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta_{t-1})-\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta^{\prime}_{t-1})\|_{F}.$
Then, according to (25), (27) and (29),
@@
which competes the proof of (16).
A.3.2 Proof of Lemma 4
According to (25),
$=$ $\Big{\|}\frac{\partial\ell(\hat{y},y_{t})}{\partial\hat{y}}\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta_{t-1})-\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta^{\prime}_{t-1})\Big{\|}_{F}$
$\leq$ $\alpha_{\ell}\Big{(}\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta_{t-1})\|_{F}+\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta^{\prime}_{t-1})\|_{F}\Big{)}$
@@ A.3.2 Proof of Lemma 4
A.4 Proof of Theorem 3
Based on Lemmas 3 and 4, we detail the proof of Theorem 3 as follows.
Note that $(\mathbf{x}_{t},y_{t})=(\mathbf{x}^{\prime}_{t},y^{\prime}_{t})$ with probability $1-\frac{1}{m}$ and $(\mathbf{x}_{t},y_{t})\neq(\mathbf{x}^{\prime}_{t},y^{\prime}_{t})$ with probability $\frac{1}{m}$. By considering (3) and incorporating the probability of the two scenarios presented in Lemmas 3 and 4, using $\mathcal{F}$ and $\mathcal{F}^{{}^{\prime}}$ to denote $f(\mathbf{x}_{t}|\theta_{t-1})$ and $f(\mathbf{x}_{t}|\theta^{\prime}_{t-1})$, respectively, we have:
@@ A.4 Proof of Theorem 3
where $\kappa_{2}=\sum_{k=1}^{K}\rho_{k}$. By (30), we have $\kappa_{2}=\nu_{\sigma}\big{(}B\alpha_{\sigma}C_{g}\big{)}^{K}C_{g}^{2}C^{2}_{\mathbf{X}}\Big{(}\sum_{j=0}^{K-1}(j+1)(B\alpha_{\sigma}C_{g})^{j}\Big{)}$, as defined in (8). Finally, since $\|\triangle\theta_{0}\|_{*}=\|\theta_{0}-\theta^{\prime}_{0}\|_{*}=0$

### v2

@@ Acknowledgment
Proof of Lemma 3.
Now, we are ready to prove Eq. (15). Firstly, note that
$\|\nabla_{\mathbf{w}}\ell(f(\mathbf{x}_{t}|\theta_{t-1}),y_{t})-\nabla_{\mathbf{w}}\ell(f(\mathbf{x}_{t}|\theta^{\prime}_{t-1}),y_{t})\|_{F}=\big\|\frac{\partial\ell(\hat{y},y_{t})}{\partial\hat{y}}\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta_{t-1})-\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta^{\prime}_{t-1})\big\|_{F}$
$\leq$ $\Big\|\Big(\frac{\partial\ell(\hat{y},y_{t})}{\partial\hat{y}}-\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}\Big)\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta_{t-1})+\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}\Big(\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta_{t-1})-\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta^{\prime}_{t-1})\Big)\Big\|_{F}$
$\leq$ $\big|\frac{\partial\ell(\hat{y},y_{t})}{\partial\hat{y}}-\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}\big|\cdot\|\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta_{t-1})\|_{F}+\big|\frac{\partial\ell(\hat{y}^{\prime},y_{t})}{\partial\hat{y}}\big|\cdot\|\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta_{t-1})-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime}_{t-1})\|_{F}$
$\leq$ $\upsilon_{\ell}\big|f(\mathbf{x}_{t}|\theta_{t-1})-f(\mathbf{x}|\theta^{\prime}_{t-1})\big|\cdot\|\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta_{t-1})\|_{F}+\alpha_{\ell}\|\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta_{t-1})-\nabla_{\mathbf{w}}f(\mathbf{x}_{t}|\theta^{\prime}_{t-1})\|_{F},$
where $\hat{y}=f(\mathbf{x}_{t}|\theta_{t-1})$ and $\hat{y}^{\prime}=f(\mathbf{x}_{t}|\theta^{\prime}_{t-1})$. Then, according to (A.7), (A.9) and (A.10), we have
@@
$\leq$ $\upsilon_{\ell}\big|f(\mathbf{x}_{t}|\theta_{t-1})-f(\mathbf{x}_{t}|\theta^{\prime}_{t-1})\big|\cdot\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}_{t}|\theta_{t-1})\|_{F}+\alpha_{\ell}\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}_{t}|\theta_{t-1})-\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}_{t}|\theta^{\prime}_{t-1})\|_{F}.$
Then, according to (A.7), (A.9) and (A.11),
@@
which competes the proof of Eq. (16).
Proof of Lemma 4.
Since $|\frac{\partial\ell(\hat{y},y)}{\partial\hat{y}}|\leq\alpha_{\ell}$ for any $\hat{y}$ and $y$, we first have that for $k=1,2,\dots,K+1$,
$\|\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}_{t}|\theta_{t-1}),y_{t})-\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}^{\prime}_{t}|\theta^{\prime}_{t-1}),y^{\prime}_{t})\|_{F}=$ $\Big\|\frac{\partial\ell(\hat{y},y_{t})}{\partial\hat{y}}\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}_{t}|\theta_{t-1})-\frac{\partial\ell(\hat{y}^{\prime},y^{\prime}_{t})}{\partial\hat{y}}\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}^{\prime}_{t}|\theta^{\prime}_{t-1})\Big\|_{F}$
$\leq$ $\alpha_{\ell}\Big(\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}_{t}|\theta_{t-1})\|_{F}+\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}^{\prime}_{t}|\theta^{\prime}_{t-1})\|_{F}\Big),$
where $\hat{y}=f(\mathbf{x}_{t}|\theta_{t-1})$ and $\hat{y}^{\prime}=f(\mathbf{x}^{\prime}_{t}|\theta^{\prime}_{t-1})$ and $\mathbf{W}^{(K+1)}=\mathbf{w}$. Finally, according to (A.7),
$\|\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}_{t}|\theta_{t-1}),y_{t})-\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}^{\prime}_{t}|\theta^{\prime}_{t-1}),y^{\prime}_{t})\|_{F}\leq$ $\alpha_{\ell}\Big(\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta_{t-1})\|_{F}+\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta^{\prime}_{t-1})\|_{F}\Big)$
@@
APPENDIX D:Proof of Theorem 3
Based on Lemma 3 and Lemma 4, we detail the proof of Theorem 3 as follows.
Note that $(\mathbf{x}_{t},y_{t})=(\mathbf{x}^{\prime}_{t},y^{\prime}_{t})$ with probability $1-\frac{1}{m}$ and $(\mathbf{x}_{t},y_{t})\neq(\mathbf{x}^{\prime}_{t},y^{\prime}_{t})$ with probability $\frac{1}{m}$. By considering Eq. (3) (in Section 3.2) and incorporating the probability of the two scenarios presented in Lemma 3 and Lemma 4, using $\mathcal{F}$ and $\mathcal{F}^{{}^{\prime}}$ to denote $f(\mathbf{x}_{t}|\theta_{t-1})$ and $f(\mathbf{x}_{t}|\theta^{\prime}_{t-1})$, respectively, we have:
@@
where $\kappa_{2}=\sum_{k=1}^{K}\rho_{k}$. By (A.12), we have $\kappa_{2}=\nu_{\sigma}\big(B\alpha_{\sigma}C_{g}\big)^{K}C_{g}^{2}C^{2}_{\mathbf{X}}\Big(\sum_{j=0}^{K-1}(j+1)(B\alpha_{\sigma}C_{g})^{j}\Big)$, as defined in (8)). Finally, since $\|\triangle\theta_{0}\|_{*}=\|\theta_{0}-\theta^{\prime}_{0}\|_{*}=0$